OpenStudy (anonymous):

Use the principle of logarithmic differentiation, where appropriate, to find the derivative of the following function.

OpenStudy (anonymous):

\[f(x)= \tan([\ln(x)]^x)\]

OpenStudy (anonymous):

let f(x) be Y \[y=\tan (\ln x)^{x}\] \[\tan^{-1} y=(lnx)^{x}\]

OpenStudy (anonymous):

\[\ln \tan^{-1} y=xln(lnx)\]

OpenStudy (anonymous):

taking derivative \[(\frac{1}{\tan^{-1} y})*(\frac{1}{1+y^{2}})*\frac{dy}{dx}=x(\frac{1}{lnx})(\frac{1}{x})+lnlnx\]

OpenStudy (anonymous):

solve for dy/dx